if a,b,c,d,e,f are in AP then (e-c) equals to
a)2(c-f)
b)2(d-b)
C)2(f-d)
D)2(d-c)

Grade:12

2 Answers

Aditya Gupta
4 years ago
hello ashish. tnx 4 d ques.
let the common difference of the AP be m.
so, c + m= d
and d+m= e
add these eqns we get
e – c= 2m
or e – c= 2(d – c)
option D
kindly approve :)
Sai Soumya
4 years ago
According to me the answer will be option d
 
FORMULA:
If a,b,c are in AP,
then 2b=a+c
Apply this , u will get the answer